Two forces 20N and 40N act simultaneously at an angle of 60º to each other. Find the magnitude

Grade:12th pass

3 Answers

Arun
3 years ago
Dear student
 
We know that resultant R² = 20² +40² + 2*20*40 cos 60
 
R² = 2000 + 800 = 2800
 
R = 20 √7 Newton
Sj
3 years ago
Hey Sereddy,Fnet=(F1^2+F2^2+2F1F2cos(a))^0.5Where F1 and F2 are the magnitude of forces and `a` is the angle between themApplying it, Fnet=20√7 NTo find direction, use parallelogram law of vectors
Vikas TU
3 years ago